OpenStudy (anonymous):

Solve the linear inequality. Other than , use interval notation to express the solution set and graph the solution set on a number line.

OpenStudy (anonymous):

\[\frac{ x }{ 6 }-\frac{ 2 }{ 3 }\le \frac{ x }{ 4 }+1\]

OpenStudy (phi):

you could multiply everything by 12 to make it \[ 2x -8 ≤ 3x+12\] now add -2x to both sides add -12 to both sides

OpenStudy (anonymous):

ok i though so.

OpenStudy (anonymous):

wait how did you get 8

OpenStudy (anonymous):

nevermind

OpenStudy (phi):

I multiplied all the terms by 12 to get rid of the fractions 12* -2/3 = -24/3 = -8

OpenStudy (anonymous):

yeah got it. thanks -20<x

OpenStudy (anonymous):

do i change the inequality sign to >

OpenStudy (phi):

isn't it -20 ≤ x use a solid circle at -20, and draw an arrow pointing to the right (toward 0)

OpenStudy (anonymous):